OpenStudy (anonymous):

Evaluate the limit s^5 - b^5 / s^2 - b^2 as s approaches b.

OpenStudy (templeguy):

−b5+s5

OpenStudy (templeguy):

−b5+s5

OpenStudy (templeguy):

-b5+5s

OpenStudy (freckles):

it is easy to see s^2-b^2=(s-b)(s+b) and the s-b on the bottom is the thing giving us problem s-b should be a factor of s^5-b^5 you can find out by divide s^5-b^5 by s-b

OpenStudy (anonymous):

well i tried googling a formula to simply a^5 - b^5 and i found the sum and difference of odd power. the formula i got was (a-b)(a^4 +(a^3)b + a^2b^2 -a(b^3) + b^4)

OpenStudy (freckles):



OpenStudy (freckles):

\[s^5-b^5=(s-b)(s^4+s^3b+s^2b^2+sb^3+b^4)\]

OpenStudy (freckles):

\[\frac{s^5-b^5}{s^2-b^2}=\frac{s^4+s^3b+s^2b^2+sb^3+b^4}{s+b}\]

OpenStudy (freckles):

here you haven gotten rid of the thing that will make the bottom zero

OpenStudy (freckles):

so now it is all about plugin' in

OpenStudy (freckles):

replacing s with b

OpenStudy (anonymous):

okay i got (5b^4)/2b

OpenStudy (freckles):

ok sounds great

OpenStudy (freckles):

I think you can reduce that :)

OpenStudy (freckles):

also our answers works when b isn't 0

OpenStudy (anonymous):

thanks!